Beam Test Meeting Mon. Apr. 16, 2012

From GlueXWiki
Jump to: navigation, search



- 3 pass (~2.5 GeV) until Thursday (photon)

- 5 pass for remainder of time after Thursday (photon)


- data taken at 5% of 2.5 GeV


- found clog in quick disconnect (Elton, Alex, Yi, John) temps dropped by few degrees C, still +3 C on downstream side, differential between chiller and downstream decreased to 1 C

- swapped flow direction from down to up to up to down (Elton, John) and downstream stayed +3C, confirming hypothesis of poor thermal connection on downstream side

- Elton will give thorough report on thermal glue used at next meeting

- data taken at different bias voltages

- still see oscillations

Future plans

- look at scalars today(FCAL)

- look at different DAQ modes tomorrow (FCAL)

- will take data at different thresholds this week (BCAL)

- install CTP in a few weeks to look at trigger options

- all detectors will remain in current location until 5 pass beam is restored

- BCAL will go to 20 degree angled position on April 25, still 23 % E